# EZIGen: Enhancing zero-shot subject-driven image generation with precise subject encoding and decoupled guidance

### v1

@@ Abstract
Zero-shot subject-driven image generation aims to produce images that incorporate a subject from a given example image. The challenge lies in preserving the subject’s identity while aligning with the text prompt, which often requires modifying certain aspects of the subject’s appearance. Despite advancements in diffusion model based methods, existing approaches still struggle to balance identity preservation with text prompt alignment. In this study, we conducted an in-depth investigation into this issue and uncovered key insights for achieving effective identity preservation while maintaining a strong balance. Our key findings include: (1) the design of the subject image encoder significantly impacts identity preservation quality, and (2) generating an initial layout is crucial for both text alignment and identity preservation. Building on these insights, we introduce a new approach called EZIGen, which employs two main strategies: a carefully crafted subject image Encoder based on the UNet architecture of the pretrained Stable Diffusion model to ensure high-quality identity transfer, following a process that decouples the guidance stages and iteratively refines the initial image layout. Through these strategies, EZIGen achieves state-of-the-art results on multiple subject-driven benchmarks with a unified model and 100 times less training data. Demo page is available at: zichengduan.github.io/pages/EZIGen/index.html.
Figure 1: Various real-life applications of EZIGen. Our model demonstrates astonishing zero-shot performances in generating high-quality images on multiple subject-driven tasks.
@@ 1 Introduction
Subject-driven generation methods allow users to create images by combining text prompts with additional subject images, adhering to the principle of ’my subject’ following ’my instructions.’ Existing solutions can be divided into two main approaches: test-time tuning-based Gal et al. (2022); Ruiz et al. (2023); Avrahami et al. (2023); Kumari et al. (2023); Li et al. (2024a); Hao et al. (2023) and zero-shot inference-based methods Wei et al. (2023); Ma et al. (2023); Purushwalkam et al. (2024); Chen et al. (2024). The former involves fine-tuning model parameters and introducing new subject tokens before generation, allowing control by combining the tuned subject token with natural language instructions. Although test-time tuning-based approaches can achieve high-quality generation, they often require a time-consuming fine-tuning step. In contrast, zero-shot approaches generate images based on the given subjects without the need for re-training, offering greater efficiency and user-friendliness. This paper focuses on advancing zero-shot methods.
Existing literature in this research area (Wei et al. (2023); Chen et al. (2024); Purushwalkam et al. (2024); Ma et al. (2023)) predominantly focuses on the efficient transfer of a referenced subject’s appearance into the generated image. These methods typically first encode the subject image and then integrate the encoded features into the UNet of the diffusion model.
@@
There are several existing options for subject image encoders. ELITE (Wei et al. (2023)) and Subject Diffusion (Ma et al. (2023)) utilize CLIP image encoders, while AnyDoor (Chen et al. (2024)) adopts DINOv2 to achieve better feature map extraction. BootPIG (Purushwalkam et al. (2024)) employs a Reference UNet model (Hu (2024)) as a subject feature extractor, reasoning that the feature space of a UNet would be more aligned with the feature space used in image generation. Although this approach is reasonable, it leaves many design aspects unexplored, such as which time step to use and how to inject the Reference UNet features into the generation UNet. Their method synchronizes the Reference UNet with the same denoising timestep as the Main UNet throughout the denoising process; however, in our study, this approach has shown sub-optimal performance Fig. 3.
@@
An often overlooked challenge in existing methods is balancing the parallel user inputs: subject image and text prompt. Although these guidances appear independent—where the subject image focuses on preserving identity and the text prompt directs the model to follow user instructions—they frequently interfere with one another. This is particularly evident when the text prompt aims to change the attribute of the subject. For instance, as shown in Fig. 3, both the text prompt “a dog in police outfit” and the subject dog image define the dog’s appearance to some extent, making it challenging to achieve a perfect balance. In such cases, prior works either prioritize identity preservation at the expense of text coherence (Ma et al. (2023)), successfully follow the text but struggle to maintain subject identity (Purushwalkam et al. (2024)), or perform suboptimally in both aspects (Wei et al. (2023); Li et al. (2024a)).
@@
In this paper, we introduce a novel subject-driven generation method called EZIGen to address the gaps mentioned earlier. Our approach innovates in both the design of the subject image encoder and the strategy for balancing identity preservation and text alignment. Building on the work of Purushwalkam et al. (2024), we utilize a Reference UNet as an extractor, reducing the complexity of aligning the feature space between the subject and generated images. However, our unique contribution lies in identifying the ‘devils in the details’: we discovered that using a fixed timestep, a frozen UNet, and coupling an adaptor for injecting identity information significantly enhances identity preservation compared to the approach used by Purushwalkam et al. (2024). To better balance subject identity preservation and text adherence, we decouple text control and subject identity injection into two distinct processes: the Layout Generation Process and the Appearance Transfer Process. The Layout Generation Process creates a coarse image layout based on the text prompt, while the Appearance Transfer Process matches the subject with the layout and transfers the extracted subject appearance features using an Adapter module, enabling explicit control over this balance. Additionally, we observe that the initial layout can impact the quality of identity injection—a layout closer to the subject tends to produce better results. Therefore, we propose an iterative pipeline, where we convert the already generated image into noisy latent space to provide a better initial layout and conduct denoising under subject guidance until a satisfactory result is reached.
@@
With the aforementioned designs, our model delivers exceptional performance in subject identity preservation while maintaining excellent text-following capabilities in subject-driven image generation tasks, offering our model fantastic abilities in generating images with various subject poses and versatile attribute modifications. Although trained on a simple synthetic dataset, our model performs remarkably well even in domains it was not specifically trained for. For instance, it can generate human facial content without dedicated pre-training. Through extensive analysis and experimentation, we demonstrate that our design surpasses previous methods across various benchmarks and tasks, achieving superior results with a unified approach.
Figure 2: Suboptimal subject image encoding. “# Sub” indicates the number of subject images used to obtain the result. Our proposed encoding preserves finer subject details while requiring fewer subject images.
Figure 3: Conflicting guidance, which previous methods often struggle to balance.
@@ 2 Related Works
enerative models are designed to synthesize samples from a data distribution based on a set of training examples. These models include Generative Adversarial Networks (GANs) (Karras et al. (2021); Goodfellow et al. (2020); Brock et al. (2018)), Variational Autoencoders (VAEs) (Kingma & Welling (2013)), autoregressive models (Esser et al. (2021); Razavi et al. (2019); Tian et al. (2024); Sun et al. (2024a)), and diffusion models (Ho et al. (2020); Song et al. (2020); Dhariwal & Nichol (2021); Rombach et al. (2022); Betker et al. (2023)). While each of these approaches has demonstrated remarkable capabilities in generating high-quality images, their inputs are typically restricted to text instructions. In comparison, our work focuses on enhancing pre-trained diffusion models by enabling them to incorporate additional image guidance alongside text prompts, providing a more comprehensive and flexible approach to image generation.
@@
Subject-driven image generation (Ruiz et al. (2023); Hao et al. (2023); Gal et al. (2022); Nam et al. (2024); Ding et al. (2024); Kumari et al. (2023); Avrahami et al. (2024); Chen et al. (2023); Liu et al. (2023)) extends the capabilities of traditional text-driven methods by incorporating additional subject images alongside text prompts, allowing for more personalized and precise image synthesis. Some methods tune a set of parameters to represent the incoming subject. For example, TextualInversion (Gal et al. (2022)) tunes a learnable text embedding to represent the subject concept, while DreamBooth (Ruiz et al. (2023)) tunes a learnable text to represent the attribute of a subject and some extra model parameters for subject class maintenance. Following this streamline, ViCo (Hao et al. (2023)) and CustomDiffusion (Kumari et al. (2023)) boost the performance and applicabilities by additionally tuning cross-attention layers. However, these models require retraining for different subjects, which limits the efficient application for productivity.
@@ 2.3 Zero-shot subject-driven image generation
To tackle the aforementioned issues, some researchers introduced zero-shot methods that accept new subjects without retraining. Some attempts like InstantBooth (Shi et al. (2024)), FastComposer (Xiao et al. (2023)), and PhotoMaker(Li et al. (2024b)) developed zero-shot generation methods only for domain-specific data, e.g. human faces. For common objects, ELITE (Wei et al. (2023)), BLIP-Diffusion (Li et al. (2024a)), Subject Diffusion (Ma et al. (2023)), and BootPIG (Purushwalkam et al. (2024)) achieved high-quality zero-shot generation by injecting detailed subject image features into diffusion spaces. However, these methods either suffer from degraded subject-identity preservation caused by sub-optimal utilization of feature extractors, or poor balancing between text and subject guidance. In this paper, we tackle these issues by deeply analyzing each problem and developing a more advanced subject-driven structure.
High-quality
subject ID?
Strong
text prompt
adherence?
Single
subject
image?
Small
training
dataset?
Can do
end-to-end
training?
Can do
multi
tasking?
  Elite
BLIP-Diffusion
BootPIG
Subject-Diffusion
Ours
Table 1: Differences between our approach and related zero-shot subject-driven methods.
@@ 3 Methods
Our method comprises two key components: an approach to encode the subject image and incorporate it into the generation process, and a strategy to balance subject identity preservation with text alignment. We will first elaborate on these two components in the context of subject-driven image generation and then discuss how our approach extends to subject-driven image editing.
Figure 4: Illustration of the proposed system. Starting from Encoding and Injecting subject feature (Sec. 3.1), where we extract a set of intermediate latent features during the simulated late denoising process of the given noisy subject image using our fixed Reference UNet, we then regard these features as offline Subject Guidance and inject it to the Main UNet via a learnable adapter. Then we showcase how we decouple the generation process into the Layout Generation Process and Appearance Transfer Process in Sec. 3.2. We first leverage the text prompt as Text Guidance using the original text-guided diffusion process to obtain a layout latent at a middle timestep $t$, then we incorporate the offline Subject Guidance to transfer the subject appearance into the layout latent in the rest of the timesteps. Finally, to achieve a complete transfer, we develop the Iterative Appearance Transfer mechanism (Sec. 3.3) to repeat the Appearance Transfer Process by adding noise to the generated image, which continues to repeat for $N$ times until satisfaction.
@@ 3.1 Encoding and Injecting Subject Image Information
The identity information of the subject image is extracted using an image encoder. Several options exist in current methods, such as CLIP, utilized in Subject Diffusion (Ma et al. (2023)), DINO-V2, as employed in AnyDoor (Chen et al. (2024)), and a Reference UNet initialized with the same parameters as Stable Diffusion, as seen in BootPIG (Purushwalkam et al. (2024)). The last option is particularly appealing because the noisy latent of a diffusion model and the encoding of the subject image are processed by similar models, potentially resulting in a closer feature space. This reduces the challenge of aligning the feature spaces of the subject image encoder and the diffusion UNet. However, this design leaves several open questions regarding how to configure the feature extractor and how to integrate the subject image encoding into the diffusion UNet. In BootPIG (Purushwalkam et al. (2024)), all parameters in the Reference UNet are open for fine-tuning, and the input image is progressively corrupted with noise at each timestep and then fed into the Reference UNet afterward, where the input subject image features are injected into the Main UNet at the same timestep. Our work identifies potential issues with this approach, as the overly-noised feature may convey inaccurate information about the subject image, leading to poor identity transfer. Therefore, in this paper, we propose an alternative solution: directly input the image with light noise into the Reference UNet to guarantee an accurate subject representation and introduce a learnable adaptor to bridge the discrepancy between latent representations from input images with different noise levels.
@@
As illustrated in Fig. 4A, we first derive a Reference UNet from the original Stable Diffusion model and fix it as an offline subject encoder, denoted as $U_{\text{ref}}(\cdot)$. We then remove the background of the subject image to discard background disturbance. Then, to initiate the last denoising step in Reference UNet and extract high-quality subject representations, we start by setting the denoising timestep $T_{\text{sub}}$ to 1 and then add noise $\epsilon$ to the subject image $x_{\text{sub}}$ accordingly:
$x_{\text{sub}}^{\prime}=\text{add\_noise}(x_{\text{sub}},\epsilon,T_{\text{sub}}),\text{where}\ T_{\text{sub}}=1$ (1)
@@
Next, we feed the slightly noised image $x_{\text{sub}}^{\prime}$, timestep $T_{\text{sub}}$ to $U_{\text{ref}}(\cdot)$ to gather the inputs of $N$ self-attention layers as a set of latent representations $[s_{1},\ s_{2}\dots s_{N}]$ of the subject image, together denoted as $\mathcal{F}_{\text{sub}}$. This procedure is formed as:
$\mathcal{F}_{\text{sub}}=[s_{1},\ s_{2}\dots s_{N}]=U_{\text{ref}}(x_{sub}^{\prime},\ T_{\text{sub}})$ (2)
@@
To inject the extracted feature, we inject an additional Adapter as an extra attention module in between the self-attention and cross-attention block within each Main UNet transformer block, denoted as $A$, resulting in $N$ Adapters in total. For each Adapter $A_{\text{n}}$, we project the Main UNet latent feature $x^{\prime}$ as queries $Q$, keys $K$, and values $V$, while using the corresponding subject feature $s_{\text{n}}$ as additional key and value tokens, denoted as $K_{\text{sub}}$ and $V_{\text{sub}}$. The output feature after each appearance transfer $\mathcal{F}_{\text{n}}$ is then obtained as follows:
$K_{\text{sub}}=\text{to\_k}(s_{\text{n}}),\ \ V_{\text{sub}}=\text{to\_v}(s_{\text{n}}),$ (3)
$\mathcal{F}_{\text{n}}=A_{\text{n}}(Q,\text{concat}[K,K_{\text{sub}}],\text{concat}[V,V_{\text{sub}}])$ (4)
where the extra keys and values are concatenated along the token dimension.
@@
3.2 Decoupling text and subject image guidance and iterative generation strategy
Figure 5: Detailed illustration of the decoupling operation. Line 1 represents the original Stable Diffusion generation process under pure text guidance. Line 2 indicates that we add subject guidance alongside text guidance in every step. Line 3 presents the decoupling operation, where we only add subject guidance in late generation steps. Line 4 shows the iterative generation design, in which we take the incomplete result from line 3 and repeatedly add noise and denoise with subject guidance until the subject identity is fully transferred.
As mentioned in the Introduction section, existing methods (Wei et al. (2023); Ma et al. (2023); Purushwalkam et al. (2024)) often struggle to balance subject identity preservation (Subject Guidance) with text adherence (Text Guidance). Although the Adaptor proposed in the previous section excels in subject identity preservation, it still faces challenges in balancing these two guidances. We observe that our proposed adaptor tends to overly prioritize subject identity preservation at the expense of text adherence (Fig. 5 line 2). This happens because the text-guided semantic layout and color patterns, such as “a teacher’s outfit”, are not fully established in the early stages and can be overshadowed by strong subject features. Therefore, instead of setting the guidances in parallel and manually adjusting the strength of each guidance using a scaling factor(Wei et al. (2023); Purushwalkam et al. (2024); Ma et al. (2023)), which often compromises one to favor the other, we decouple each guidance to allow them to dominate at different stages: text-guidance occurs in early stages, while fine-grained appearance details emerge later. Consequently, this results in two sub-processes: namely the Layout Generation Process and the Appearance Transfer Process (Fig. 5 line 3).
Layout Generation Process. First, we take the original text prompt as Text Guidance and generate a coarse layout using the Stable Diffusion model. Specifically, we interrupt the generation process at a certain timestep $T_{\text{layout}}$ and regard the intermediate latent as the coarse layout latent, denoted as $x^{\prime}$, containing the overall semantic structure and rough color patterns of the image, as shown by the last image of the orange box in Fig. 5 line 3.
Appearance Transfer Process. Then, as shown in the blue box in Fig. 5 line 4, we bring in subject feature $\mathcal{F}_{\text{sub}}$ as Subject Guidance and transfer the subject appearance to the layout $x^{\prime}$ using the adapters $A$. Intuitively, the attention mechanism within the adaptor first matches (represented by the attention map) subject image patches from the Reference UNet to the noisy latents from the Main generation UNet, which define the scene’s initial layout. It then transfers the content (encoded by the V values) from the subject patches to their corresponding locations in the image being generated. This understanding can be illustrated in Fig. 4, given a rough layout from the Layout Generation Process depicting “a dog in a teacher outfit”, our model will first match between the subject dog and the dog in the rough latent layout, then transfer only the paired brown furry skin and maintain the blue teacher’s outfit untouched as it is not matched with the dog’s fur.

### v2

@@ Abstract
Zero-shot subject-driven image generation aims to produce images that incorporate a subject from a given example image. The challenge lies in preserving the subject’s identity while aligning with the text prompt which often requires modifying certain aspects of the subject’s appearance. Despite advancements in diffusion model based methods, existing approaches still struggle to balance identity preservation with text prompt alignment. In this study, we conducted an in-depth investigation into this issue and uncovered key insights for achieving effective identity preservation while maintaining a strong balance. Our key findings include: (1) the design of the subject image encoder significantly impacts identity preservation quality, and (2) separating text and subject guidance is crucial for both text alignment and identity preservation. Building on these insights, we introduce a new approach called EZIGen, which employs two main strategies: a carefully crafted subject image Encoder based on the pretrained UNet of the Stable Diffusion model to ensure high-quality identity transfer, following a process that decouples the guidance stages and iteratively refines the initial image layout. Through these strategies, EZIGen achieves state-of-the-art results on multiple subject-driven benchmarks with a unified model and 100 times less training data. The demo page is available at: https://zichengduan.github.io/pages/EZIGen/index.html.
Figure 1: Our model demonstrates remarkable zero-shot performances in generating high-quality images on subject-driven generation and editing tasks.
@@ 1 Introduction
Figure 2: Suboptimal encoding. BootPIG’s encoder design may lead to suboptimal performance compared to our own. “Ours” means replacing BootPIG’s encoder with our encoder design.
Figure 3: Conflicting guidance. Existing methods struggle to strike a good balance between identity preservation and text prompt alignment.
@@
Subject-driven generation methods enable users to create images by combining text prompts with subject images, following the principle of ’my subject’ following ’my instructions.’ Existing solutions fall into two categories: test-time tuning-based(Gal et al. (2022); Ruiz et al. (2023); Avrahami et al. (2023); Kumari et al. (2023); Li et al. (2024a); Hao et al. (2023)) and zero-shot inference-based(Wei et al. (2023); Ma et al. (2023); Purushwalkam et al. (2024); Chen et al. (2024)). Test-time tuning involves fine-tuning model parameters and introducing subject tokens, allowing detailed control but requiring time-consuming re-training. In contrast, zero-shot methods generate images directly from the given subjects without re-training, offering greater efficiency. This paper focuses on improving zero-shot methods.
@@
Existing zero-shot methods (Wei et al. (2023); Chen et al. (2024); Purushwalkam et al. (2024); Ma et al. (2023)) predominantly focus on the transfer of a reference subject’s appearance into the generated image. These methods typically first encode the subject image and then integrate the encoded features into the UNet of the diffusion model. There are several existing options for subject image encoders. ELITE (Wei et al. (2023)) and Subject Diffusion (Ma et al. (2023)) utilize CLIP image encoders, while AnyDoor (Chen et al. (2024)) adopts DINOv2 to achieve better feature map extraction. BootPIG (Purushwalkam et al. (2024)) employs a Reference UNet model (Hu (2024)) as a subject feature extractor, reasoning that the feature space of a UNet would be more aligned with the feature space used in image generation. Although this approach is reasonable, it leaves many design aspects unexplored, such as which time step to use and how to inject the Reference UNet features into the generation UNet. Our research reveals that those aspects might have a big impact at the identity preservation capability.
@@
An often overlooked challenge in existing methods is balancing the parallel user inputs: subject image and text prompt. Although these guidances appear independent—where the subject image focuses on preserving identity and the text prompt directs the model to follow user instructions—they frequently interfere with one another. For instance, as shown in Fig. 3, both the text prompt “a dog in police outfit” and the subject dog image define the dog’s appearance to some extent simultaneously. In such cases, prior works either prioritize identity preservation at the expense of text coherence (Ma et al. (2023)), successfully follow the text but struggle to maintain subject identity (Purushwalkam et al. (2024)), or perform suboptimally in both aspects (Wei et al. (2023); Li et al. (2024a)).
@@
In this paper, we present a novel subject-driven generation method, EZIGen, addressing the challenges of identity preservation and text alignment. Building on Purushwalkam et al. (2024), we employ a Reference UNet as an extractor to achieve good feature alignment between subject and generated images. However, our unique contribution lies in identifying the ‘devils in the details’: we discovered that using a fixed timestep, a frozen UNet, and coupling an adaptor for injecting identity information significantly enhances identity preservation compared to the approach used by BootPIG Purushwalkam et al. (2024). To better balance subject identity and text adherence, we decouple the generation process into two distinct stages: the Layout Generation Process, which forms a coarse layout from text prompts, and the Appearance Transfer Process, which injects the encoded subject details via the adapter. This decoupling explicitly separates guidance signals. Additionally, we observe that the initial layout can impact the quality of identity injection—a layout closer to the subject tends to produce better results. Therefore, we introduce an iterative pipeline that converts the generated image back into an editable noisy latent, refining the layout with subject guidance.
@@
With the aforementioned designs, our model delivers exceptional performance in subject identity preservation while maintaining excellent text-following capabilities in subject-driven image generation tasks, offering outstanding abilities in generating images with various subject poses and versatile attribute modifications. Furthermore, our model consistently performs remarkably well even in domains for which it was not specifically trained or fine-tuned. For instance, it can effortlessly generate highly detailed human facial content without any dedicated pre-training or domain-specific adjustments. Through extensive analysis, rigorous testing, and comprehensive experiments, we demonstrate that our design consistently surpasses previous methods across various benchmarks and tasks, achieving superior and reliable results with a unified, efficient approach.
@@ 2 Related Works
Generative models are designed to synthesize samples from a data distribution based on a set of training examples. These models include Generative Adversarial Networks (GANs) (Karras et al. (2021); Goodfellow et al. (2020); Brock et al. (2018)), Variational Autoencoders (VAEs) (Kingma & Welling (2013)), autoregressive models (Esser et al. (2021); Razavi et al. (2019); Tian et al. (2024); Sun et al. (2024a)), and diffusion models (Ho et al. (2020); Song et al. (2020); Dhariwal & Nichol (2021); Betker et al. (2023)). While each of these approaches has demonstrated remarkable capabilities in generating high-quality and diverse images, their inputs are typically restricted to text instructions or predefined conditions. In contrast, our work significantly enhances pre-trained diffusion models by enabling them to incorporate additional image guidance alongside text prompts, ultimately providing a more comprehensive, versatile, and flexible approach to image generation across a wider range of applications and contexts.
@@
Tuning-based subject-driven image generation (Ruiz et al. (2023); Hao et al. (2023); Gal et al. (2022); Nam et al. (2024); Ding et al. (2024); Kumari et al. (2023); Avrahami et al. (2024); Chen et al. (2023); Liu et al. (2023)) typically adjusts sets of parameters to extend traditional text-driven methods, allowing them to incorporate additional subject images alongside text prompts, thereby enabling more personalized, detailed, and flexible image synthesis. Some approaches focus on tuning text embeddings to represent the subject accurately, such as TextualInversion (Gal et al. (2022)), which simply adjusts a learnable text embedding, and DreamBooth (Ruiz et al. (2023)), which fine-tunes both text embeddings and model parameters for more precise and effective control. ViCo (Hao et al. (2023)) and CustomDiffusion (Kumari et al. (2023)) further improve performance by additionally tuning cross-attention layers, enhancing subject appearance integration. Despite these advancements, these methods require re-training for each individual subject, making them time-consuming, and unsuitable for productivity in large-scale applications or practical deployments.
@@ 2.3 Zero-shot subject-driven image generation
To tackle the aforementioned issues, some researchers introduced zero-shot methods that accept new subjects without requiring retraining. Some attempts, such as InstantBooth (Shi et al. (2024)), FastComposer (Xiao et al. (2023)), and PhotoMaker (Li et al. (2024b)), developed zero-shot generation techniques specifically for domain-specific data, such as human faces or other constrained applications. For more general objects, ELITE (Wei et al. (2023)), BLIP-Diffusion (Li et al. (2024a)), Subject Diffusion (Ma et al. (2023)), and BootPIG (Purushwalkam et al. (2024)) achieved high-quality zero-shot generation by injecting detailed subject image features into diffusion spaces. Nevertheless, these methods still struggle with issues like degraded subject-identity preservation due to sub-optimal feature extractor utilization or poor balancing between the text and subject guidance during the generation process.
@@ 3 Methods
Our method comprises two main components: a technique to encode the subject image for the generation process and a strategy to balance subject identity preservation with text alignment. We will first elaborate on these components for subject-driven image generation and then extend the discussion to subject-driven image editing.
Figure 4: Illustration of the proposed system. We begin by Encoding and Injecting subject features (Sec. 3.1). Next, we decouple the generation into the Layout Generation Process and Appearance Transfer Process (Sec. 3.2). Finally, we introduce the Iterative Appearance Transfer mechanism (Sec. 3.3) to fully transfer the subject appearance feature to the layout image.
@@ 3.1 Encoding and Injecting Subject Image Information
The identity information of the subject image is extracted using an image encoder. Several options exist in current methods, such as CLIP, utilized in Subject Diffusion (Ma et al. (2023)), DINO-V2, as employed in AnyDoor (Chen et al. (2024)), and a Reference UNet initialized with the same parameters as Stable Diffusion, as seen in BootPIG (Purushwalkam et al. (2024)). The last option is particularly appealing because the noisy latent of a diffusion model and the encoding of the subject image are processed by similar models, potentially resulting in a closer feature space. This reduces the challenge of aligning the feature spaces of the subject image encoder and the diffusion UNet. However, this design leaves several open questions regarding how to configure the feature extractor and how to integrate the subject image encoding into the diffusion UNet. In BootPIG (Purushwalkam et al. (2024)), all parameters in the Reference UNet are open for fine-tuning, and the input image is progressively corrupted with noise at each timestep and then fed into the Reference UNet afterward, where the input subject image features are injected into the Main UNet at the same timestep. Our work identifies potential issues with this approach: the overly-noised feature may convey inaccurate information about the subject image, and tunning the Reference UNet would disrupt the Stable Diffusion parameters, leading to suboptimal subject encoding(Fig. 2 shows some examples). In this paper, we propose an alternative solution: inputting the image with light noise into a fixed Reference UNet to ensure accurate subject representation and introducing a learnable adapter to bridge discrepancies between latent representations from images with varying noise levels. As shown in Table 5 and Figure 9, our approach significantly improves identity preservation compared to existing methods.
@@
As depicted in the gray box in Fig. 4, we derive a Reference UNet from the original Stable Diffusion model and employ it as a fixed offline subject encoder, denoted as $U_{\text{ref}}(\cdot)$. To minimize background interference, the subject image’s background is removed. For extracting subject representations, we set the denoising timestep $T_{\text{sub}}=1$ and add Gaussian noise $\epsilon$ to the subject image $x_{\text{sub}}$, depending on the $T_{\text{sub}}$ and a noise scheduler $\phi$:
$x_{\text{sub}}^{\prime}=x_{\text{sub}}+\phi(\epsilon,T_{\text{sub}}),\quad\text{where}\ \epsilon\sim\mathcal{N}(0,\sigma^{2}),\quad T_{\text{sub}}=1$ (1)
@@
This slightly noised image $x_{\text{sub}}^{\prime}$ and the timestep $T_{\text{sub}}$ are then passed through $U_{\text{ref}}(\cdot)$ to obtain the latent representations from all $N$ self-attention layers. These representations, collectively denoted as $\mathcal{F}_{\text{sub}}$, capture the subject-specific features:
$\mathcal{F}_{\text{sub}}=\{s_{1},s_{2},\ldots,s_{N}\}=U_{\text{ref}}(x_{\text{sub}}^{\prime},T_{\text{sub}})$ (2)
@@
To integrate these subject features, we introduce an Adapter as an additional attention module situated between the self-attention and cross-attention blocks within each transformer block of the Main UNet, resulting in a total of $N$ Adapters. For each Adapter $A_{n}$, the Main UNet’s latent feature $x^{\prime}$ is projected into query, key, and value matrices $Q$, $K$, and $V$, respectively. Meanwhile, the subject feature $s_{n}$ is mapped into additional key $K_{\text{sub}}$ and value $V_{\text{sub}}$ matrices as follows:
$K_{\text{sub}}=W_{k}s_{n},\quad V_{\text{sub}}=W_{v}s_{n}$ (3)
The output feature $\mathcal{F}_{n}$ for each Adapter $A_{n}$ is then computed by combining the Main UNet’s features with the subject features through the following attention operation:
$\mathcal{F}_{n}=A_{n}(Q,[K;K_{\text{sub}}],[V;V_{\text{sub}}])$ (4)
Here, we have $\forall n\in\{1,2,\ldots,N\}$, $[K;K_{\text{sub}}]$ and $[V;V_{\text{sub}}]$ represent the concatenation of feature vector $K$ with $K_{\text{sub}}$ and $V$ with $V_{\text{sub}}$ along the token dimension.
@@
3.2 Decoupling text and subject guidance
As mentioned previously, existing methods (Wei et al. (2023); Ma et al. (2023); Purushwalkam et al. (2024)) often struggle to balance subject ID preservation, i.e. subject guidance, with text adherence, i.e. text guidance. While the design in Sec. 3.1 excels in preserving subject identity, it still faces challenges in achieving this balance. We observe that injecting subject features alongside text prompts at all timesteps tends to prioritize subject identity, overshadowing text-guided semantic layouts and color patterns that are not fully established in the early stages, such as the “teacher’s outfit.” Instead of using parallel guidance with scaling factors (Wei et al. (2023); Purushwalkam et al. (2024); Ma et al. (2023)), which often compromises one aspect, we decouple the guidances to let them dominate at different stages: text guidance in the early stages and subject guidance details later. This leads to two distinct sub-processes: the Layout Generation and Appearance Transfer Process.
Layout Generation Process. First, we take the original text prompt as text guidance and generate a coarse layout using the Stable Diffusion model. Specifically, we interrupt the generation process at a certain timestep $T_{\text{layout}}$ and regard the intermediate latent as the coarse layout latent, denoted as $x^{\prime}$, containing the overall semantic structure and rough color patterns of the image, as shown by the last image of the orange box in Fig. 4.
Appearance Transfer Process. Then, as shown in the blue box in Fig. 4, we bring in subject feature $\mathcal{F}_{\text{sub}}$ as subject guidance and transfer the subject appearance to the layout $x^{\prime}$ using the adapters $A$. Intuitively, we discover that the attention mechanism within the adaptor first establishes matching (represented by the attention map) between subject image patches from the Reference UNet and the noisy latents from the Main generation UNet, which define the scene’s initial layout. It then transfers the content (encoded by the V values) from the subject patches to their corresponding locations in the image being generated. This understanding can be illustrated in Fig. 4, given a rough layout from the Layout Generation Process depicting “a dog in a teacher outfit”, our model will first match between the subject dog and the dog in the rough latent layout, then transfer only the paired brown furry skin and maintain the blue teacher’s outfit untouched.